McGown Corp has the following information:

Beginning Inventory (1/1) Ending Inventory (12/31)
Raw Materials Inventory $ 24,900 $ 33,400
Work in Process Inventory $ 22,700 $ 23,300
Finished Goods Inventory $ 33,200 $ 23,800


Additional information for the year is as follows:

Raw materials purchases $ 101,200
Direct labor $ 83,300
Manufacturing overhead applied $ 89,100
Indirect materials $ 0


Compute the unadjusted cost of goods sold.

$273,900

$264,500

$274,500

$147,200
Answer #1

Direct material used = 24900+101200-33400 = 92700

Total manufacturing cost = 92700+83300+89100 = 265100

Cost of goods manufactured = 22700+265100-23300 = 264500

Unadjusted cost of goods sold = 33200+264500-23800 = 273900

So answer is a) $273900